OpenStudy (anonymous):

Wendy opens a bank account with money she has saved. After this initial deposit, Wendy then deposits the same amount of money each week. Using the table below, how much was her initial deposit? x y Week Balance in Dollars 1 76 2 92 3 108 4 124 A. $60 B. $76 C. $16 D. $50

OpenStudy (anonymous):

x is 1, 2,3, 4 and the y is the bigger numbers

OpenStudy (anonymous):

i need the answer someone help!!!

OpenStudy (anonymous):

Tom starts a lawnmowing business. Before he can begin, he purchases a new lawnmower. Each weekend he mows the same number of lawns and earns the same amount of money. Write an equation to show how much money he's earned at any week. x y Week Profit in Dollars 1 –90 2 –30 3 30 4 90 A. y = 60x – 90 B. y = 60x – 150 C. y = 90x + 150 D. y = 60x

OpenStudy (xguardians):

The First question is A) 60$ Why? The Rate of change between them is $16 and then you subtract that from 76$.

OpenStudy (xguardians):

Second question is B) y = 60x – 150 Why? From the chart you can see that he makes 60$ each time, therefore the x value must be 60. Since after one week he gains 60, you subtract 60 from 90 to get the initial start which was -150.

OpenStudy (anonymous):